An MTN cellular-phone user notices that her cellular phone accounts for three
consecutive months were significantly different. In May, her account was one and a half
times more than her account in June. Her account in July was R 50 more than her account in
June. In total, she spent R575 in cellular-phone accounts in the three months. What was her
account for each month?
1
Expert's answer
2019-04-28T18:01:44-0400

Let "x" be the money she spent in June, then in July she spent "(x+50)", and in May she spent "1.5\times x". As in total she spent "575", then

"1.5\times x +(x+50)+x=575"

"3.5\times x = 575-50"

"x=150"

Then in June she spent "150", in July "150+50=200", and in May"150\times 1.5= 225"